# secReferences

### v1

@@ Abstract
Knowledge distillation (KD) has been widely studied in unsupervised Industrial Image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information. In this paper, we propose Crossmodal Reverse Distillation (CRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on the MVTec 3D-AD dataset demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code is available at CRD.
@@ 1 Introduction
Figure 1: Comparison of our proposed Crossmodal Reverse Distillation (CRD) with other multimodal AD methods based on feature learning. The upper part contains an overview of the two previous categories of methods, and the lower part shows the overall framework of CRD.
Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [18] and M3DM [28]. These methods construct large memory banks by processing and storing the multimodal features of normal samples from the training set. During inference, the extracted features are compared with the normal features in the memory bank to determine if the samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
@@
Currently, feature learning-based methods are divided into two categories as in Fig. 1: (1) Methods like AST [24] and MMRD [15] (Fig. 1 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples. However, these methods may smooth out abnormal features during fusion, reducing teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous. (2) Crossmodal Feature Mapping [12] (Fig. 1 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. But inevitably, misalignment between multimodal features results in underfitting and cause false positives in the output anomaly maps.
@@
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD), named Crossmodal Reverse Distillation (CRD). CRD naturally combines the strengths of both feature learning paradigms. First, we design a Multi-branch Distillation (MBD) that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two crossmodal assistants within each modality branch. One is Crossmodal Filter (CF), which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier (CA), which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of the test samples are calculated by integrating the anomaly maps of all branches. Experimental results on the representative 3D AD dataset demonstrate the effectiveness of our method in multimodal anomaly detection and localization. Our contributions are as follows:
  • We propose Crossmodal Reverse Distillation based on Multi-branch Distillation for multimodal industrial anomaly detection, which effectively detects anomalies in all modalities.
  • Two crossmodal assistants are introduced including Crossmodal Filter that helps the student decoder generate normal features, and Crossmodal Amplifier that amplifies the perception of anomalies from other modalities.
  • Experimental results show that our proposed CRD achieves state-of-the-art performance on the multimodal industrial AD dataset.
@@ 2 Related Work
Unsupervised Anomaly Detection
Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [5, 1, 34]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [23, 2, 20], comparing test samples with stored normal features to detect anomalies. Recently, there has been significant progress in the research of synthetic anomaly images [19, 30, 21], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [25, 27, 24, 3, 13, 26, 16, 14, 33, 17] based on teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
Multimodal Anomaly Detection
With the release of multiple 3D industrial anomaly detection datasets [7, 9, 22], unsupervised multimodal AD has gradually become a topic of research. Some methods [11], such as BTF [18] and M3DM [28], follow PatchCore [23] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [10, 31, 32] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [4], AST [24], and MMRD [15], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two lightweight neural networks to perform crossmodal feature mapping and locate anomalies based on the mapping results.
@@ 3 Preliminaries
$\mathrm{Sim}(f_{1},f_{2})=\frac{{f_{1}}^{\mathsf{T}}f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
@@
where $l$ represents the number of selected feature layers, typically set to 3. During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map, and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample. When the input is a normal sample, the student’s features are similar to the teacher’s, thus both the values in $M$ and $S$ are low. When the input is an anomalous sample, since the student has not learned the teacher’s abnormal feature representation ability, the feature distance $M$ is large in the anomalous regions, allowing for the detection of both the presence and location of anomalies.
Reverse distillation (RD) [13] is an advanced KD-based AD method. Unlike traditional forward distillation, RD uses an encoder-decoder architecture. The teacher network is a pre-trained encoder with a WideResNet 50 [29] backbone, while the student network is a trainable decoder that generates features matching the size of the teacher’s features. RD also includes a trainable bottleneck module, OCBE, to compress features between the teacher and student. During training, the student network learns to reconstruct the teacher’s normal features using normal samples. During inference, the teacher network, with its strong feature extraction ability, captures abnormal patterns in the anomalous input and generates anomaly features. At the same time, the OCBE module compresses the teacher features, preventing abnormal information from inputting student network. As a result, even for abnormal samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
@@ 4 Crossmodal Reverse Distillation
Figure 2: Pipeline of our proposed method which introduces Knowledge Distillation to the multimodal industrial AD task. Our proposed method mainly consists of two parts: (1) Cross Reverse Distillation designs multiple branches to realize Reverse Distillation for multiple modalities, (2) Anomaly Scoring aggregates anomaly maps from multiple branches to achieve multimodal anomaly detection.
@@
In the multimodal anomaly detection task, each input consists of multiple modalities of data. In this section, we primarily focus on RGB and Depth modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{RGB},\mathcal{I}^{Dep}\}$, where $\mathcal{I}^{RGB}$ and $\mathcal{I}^{Dep}$ are the RGB image and depth map respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which includes both normal and anomalous samples.
@@ 4.1 Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories:
  • Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem.
  • Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network.
The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features from the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
@@
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 2, which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including RGB and depth branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality, eliminating the need for fusion at the input or feature exaction stages. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
For the design of each branch, we draw inspiration from RD. Both the RGB and depth branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec. 4.2). The teacher encoders in both branches share parameters and are based on WideResNet 50 pre-trained on ImageNet. The student decoders follow the design of RD, using a structure symmetric to that of the teacher networks. Crossmodal Filter replaces the OCBE module of RD, while Crossmodal Amplifier is placed after the student decoder to optimize the output features, enabling crossmodal mapping and information fusion between modalities. The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
DepthGAN [7] 53.8 37.2 58.0 60.3 43.0 53.4 64.2 60.1 44.3 57.7 53.2
DepthAE [7] 64.8 50.2 65.0 48.8 80.5 52.2 71.2 52.9 54.0 55.2 59.5
DepthVM [7] 51.3 55.1 47.7 58.1 61.7 71.6 45.0 42.1 59.8 62.3 55.5
BTF [18] 91.8 74.8 96.7 88.3 93.2 58.2 89.6 91.2 92.1 88.6 86.5
M3DM [28] 99.4 90.9 97.2 97.6 96.0 94.2 97.3 89.9 97.2 85.0 94.5
Shape-guided [11] 98.6 89.4 98.3 99.1 97.6 85.7 99.0 96.5 96.0 86.9 94.7
AST [24] 98.3 87.3 97.6 97.1 93.2 88.5 97.4 98.1 100 79.7 93.7
MMRD [15] 99.9 94.3 96.4 94.3 99.2 91.2 94.9 90.1 99.4 90.1 95.0
CFM [12] 98.8 87.5 98.4 99.2 99.7 92.4 96.4 94.9 97.9 95.0 96.0
CRD (Ours) 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
Table 1: Image-level anomaly detection results I-AUROC (%) on MVTec 3D-AD with the best in bold. We abbreviate Crossmodal Feature Mapping as CFM, the same below.
@@ 4.2 Crossmodal Assistants
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two crossmodal assistants serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final detection.
Crossmodal Filter
The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
@@
Crossmodal Amplifier
When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec. 4.1. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
@@
In contrast to Bottleneck Projection in CF, the proposed Inverted Bottleneck Projection first expands the dimensions of features and then compresses them to the original sizes. We believe that this process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information in the modality during feature mapping.

### v2

# Multimodal Industrial Anomaly Detection by Crossmodal Reverse Distillation

@@ Abstract
Knowledge distillation (KD) has been widely studied in unsupervised image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information. In this paper, we propose Crossmodal Reverse Distillation (CRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code will be available.
@@ 1 Introduction
Figure 1: Qualitative results on MVTec 3D-AD [7]. From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [26] with more false negatives, Crossmodal Feature Mapping [12] (abbreviated as CFM) with more false positives, and our method CRD with superior segmentation.
Figure 2: Comparison of our proposed CRD with other multimodal AD methods based on feature learning.
Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [20] and M3DM [31]. These methods construct large memory banks by processing and storing multimodal features of normal samples. During inference, extracted features are compared with normal features in the memory bank to determine if these samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
@@
Currently, feature learning-based methods are divided into two categories as in Fig. 2: (1) Methods like AST [26] and MMRD [15] (Fig. 2 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples. However, these methods may smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous. (2) Crossmodal Feature Mapping [12] (Fig. 2 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
@@
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [13], named Crossmodal Reverse Distillation (CRD). CRD naturally combines the strengths of both feature learning paradigms. First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two crossmodal assistants within each modality branch. One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier, which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches. Experimental results on representative multimodal AD datasets demonstrate the effectiveness of CRD in multimodal anomaly detection and localization.
@@ 2 Related Work
Unsupervised Anomaly Detection. Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [5, 1, 38, 32, 19, 18]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [25, 2, 22], comparing test samples with stored normal features to detect anomalies. Recently, significant progress has been made in the research of synthetic anomaly images [21, 34, 23], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [27, 30, 26, 3, 13, 28, 16, 14, 37, 17] based on the teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
Multimodal Anomaly Detection. With the release of multiple 3D industrial anomaly detection datasets [7, 9, 24], unsupervised multimodal AD has gradually become a topic of research. Some methods [11, 29], such as BTF [20] and M3DM [31], follow PatchCore [25] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [10, 35, 36] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [4], AST [26], and MMRD [15], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
@@ 3 Preliminaries
$\mathrm{Sim}(f_{1},f_{2})=\frac{f_{1}\cdot f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
@@
where $l$ represents the number of selected feature layers, typically set to 3. During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map. , and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample. When the input is a normal sample, the student’s features are similar to the teacher’s, thus both the values in $M$ and $S$ are low. When the input is an anomalous sample, since the student has not learned the teacher’s abnormal feature representation ability, the feature distance $M$ is large in the anomalous regions, allowing for the detection of both the presence and location of anomalies.
Reverse distillation (RD) [13] is an advanced KD-based AD method. Unlike traditional forward distillation, RD uses an encoder-decoder architecture. The teacher network is a pre-trained encoder WideResNet-50 [33], while the student network is a trainable decoder that generates features matching the size of the teacher’s features. RD also includes a trainable one-class bottleneck embedding (OCBE) module to compress the teacher’s features brefore passing them to the student. During training, the student network learns to reconstruct the teacher’s normal features using normal samples. During inference, the teacher network, with its strong feature extraction ability, captures abnormal patterns in the anomalous input and generates anomaly features. At the same time, the OCBE module compresses the teacher features, preventing abnormal information from inputting student network. Hence, even for anomalous samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
@@ 4 Crossmodal Reverse Distillation
Figure 3: Pipeline of our proposed method. The feature evolution visualization highlights that anomalies in each modality are effectively captured. In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
@@
In the multimodal AD task, each input consists of multiple modalities of data. In this section, we primarily focus on 2D (e.g., RGB images) and 3D (e.g., depth maps) modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{2D},\mathcal{I}^{3D}\}$, where $\mathcal{I}^{2D}$ and $\mathcal{I}^{3D}$ are the 2D and 3D modalities respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which includes both normal and anomalous samples.
@@ 4.1 Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories: (1) Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem [26]. (2) Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network [15]. The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features from the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
@@
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 3, which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
For the design of each branch, we draw inspiration from RD. Both 2D and 3D branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec. 4.2). The teacher encoders in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet. The student decoders follow RD, using a structure symmetric to that of the teacher networks. Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to optimize the output features, enabling crossmodal mapping and information fusion between modalities. The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
@@ 4.2 Crossmodal Assistants
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two crossmodal assistants serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
Crossmodal Filter. The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
@@
Crossmodal Amplifier. When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec. 4.1. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
@@
In contrast to Bottleneck Projection in CF, the proposed Inverted Bottleneck Projection first expands the dimensions of features and then compresses them to the original sizes. This process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information during feature mapping.
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Avg.
DepthGAN [7] 53.8/42.1 37.2/42.2 58.0/77.8 60.3/69.6 43.0/49.4 53.4/25.2 64.2/28.5 60.1/36.2 44.3/40.2 57.7/63.1 53.2/47.4
DepthAE [7] 64.8/43.2 50.2/15.8 65.0/80.8 48.8/49.1 80.5/84.1 52.2/40.6 71.2/26.2 52.9/21.6 54.0/71.6 55.2/47.8 59.5/48.1
DepthVM [7] 51.3/38.8 55.1/32.1 47.7/19.4 58.1/57.0 61.7/40.8 71.6/28.2 45.0/24.4 42.1/34.9 59.8/26.8 62.3/33.1 55.5/33.5
BTF [20] 91.8/97.6 74.8/96.9 96.7/97.9 88.3/97.3 93.2/93.3 58.2/88.8 89.6/97.5 91.2/98.1 92.1/95.0 88.6/97.1 86.5/95.9
M3DM [31] 99.4/97.0 90.9/97.1 97.2/97.9 97.6/95.0 96.0/94.1 94.2/93.2 97.3/97.7 89.9/97.1 97.2/97.1 85.0/97.5 94.5/96.4
Shape-guided [11] 98.6/98.1 89.4/97.3 98.3/98.2 99.1/97.1 97.6/96.2 85.7/97.8 99.0/98.1 96.5/98.3 96.0/97.4 86.9/97.5 94.7/97.6
AST [26] 98.3/97.0 87.3/94.7 97.6/98.1 97.1/93.9 93.2/91.3 88.5/90.6 97.4/97.9 98.1/98.2 100/88.9 79.7/94.0 93.7/94.4
MMRD [15] 99.9/98.6 94.3/99.0 96.4/99.1 94.3/95.1 99.2/99.0 91.2/90.1 94.9/99.0 90.1/99.0 99.4/98.7 90.1/98.2 95.0/97.6
CFM [12] 98.8/98.0 87.5/96.6 98.4/98.2 99.2/94.7 99.7/95.9 92.4/96.7 96.4/98.2 94.9/98.3 97.9/97.6 95.0/98.2 96.0/97.2
CRD (Ours) 99.6/97.7 96.1/98.1 98.2/99.1 98.5/98.2 99.7/98.7 90.1/91.0 99.2/99.4 97.6/99.5 100/98.2 83.3/97.4 96.2/97.7
Table 1: Anomaly detection and localization results I-AUROC/PRO (%) on MVTec 3D-AD with the best in bold and the second-best underlined.